OpenStudy (anonymous):

[Part 1 (4 points)] Show all work in dividing 4t^2-1/15t+10 divided 4t^2-4t+1/3t^2+5t+2 . [Part 2 (2 points)] Use complete sentences to explain your process.

OpenStudy (anonymous):

Is this the way it looks?\[\frac{\frac{4t ^{2}-1}{15t+10}}{\frac{4t ^{2}-4t+1}{3t ^{2}+5t+2}}\]

OpenStudy (anonymous):

Factor all 4 polynomials: The top is a difference of squares. The next one down is greatest common factor using 5. I would use ac method on the bottom fraction's trinomials. That should get you started... I'm gonna go work it.

OpenStudy (anonymous):

For this one, it seems easier to me to invert and multiply the bottom quotient. Then cancel any common factors between the numerator and denominator. My instructor always encouraged us to find a lowest common denominator, multiply the LCD by the numerator and denominator, and then reduce common factors between numerator and denominator. Use the method that your instructor has encouraged.

OpenStudy (anonymous):

Thank you so much!! That was a huge help! I really appreciate it!